Given that cosA= -3/8, the one true statement is?
1. tan A = square root55/3, only
2. sin A= + with minus underneath square root 55/8
3. sin A= + with minus sugn under 3 square root2/4
4. sec A= + with minus under 8/3

Answer by eperette(173)

cosA=-3/8 ....this means that your angle is either on quadrant II or III....
Know: you have a point (-3, y) and radius (r=8)
x^2 + y^2 = r^2
(-3)^2 + y^2 = (8)^2
9 + y^2 = 64
y^2 = 55
y = sqr(55) or y= -sqr(55) where sqr means square root
Remember:
cosA=x/r=-3/8
sinA=y/r= +- (sqr(55)/8)
tanA=y/x= +- (sqr(55)/-3)
secA=r/x = 8/-3
cscA=r/y = 8/+-(sqr(55))
cotA=x/y= -3/+-(sqr(55))
It seems like 2 would be right based on your description